OpenStudy (anonymous):

a person is considered to have a certain symptom when his or her temperature is higher than 100.94 F. The formula is F=9/5c+32 can used to convert Celsius temperatures c to Fahrenheit temperature having trouble solving the problem

OpenStudy (skullpatrol):

Use the idea of balance to solve the equation for F.

OpenStudy (anonymous):

the formula that I gave... but im lost at how to start...

OpenStudy (skullpatrol):

$$\Huge F=\frac{9}{5} C+32$$

OpenStudy (skullpatrol):

So now you want C = ???

OpenStudy (skullpatrol):

Subtract 32 from both sides please

OpenStudy (skullpatrol):

$$\Huge F-32 = \frac{9}{5} C$$

OpenStudy (skullpatrol):

now multiply both sides by 5/9

OpenStudy (skullpatrol):

$$\Huge \frac{5}{9}(F-32) = C$$

OpenStudy (skullpatrol):

now substitute 100.94 for F and work out what C will be in degrees Celsius

OpenStudy (skullpatrol):

https://en.wikipedia.org/wiki/Celcius